# Eagle: Exploring The Design Space for Multimodal LLMs with Mixture of Encoders

### v1

@@ 3.2 Main results
4 Related work
4.1 Multimodal large language models
Our work is related to the general architecture design of multimodal large language models. Besides the line of representative open-source research mentioned in the introduction section, other notable families of MLLMs include, but are not limited to MiniGPT-4 [85, 7], Lynx [82], Otter [40, 39], Qwen-VL [5], CogVLM [75, 29], VILA [48], GPT-4V [2], Gemini [71], and Llama 3.1 [19]. Depending on how vision signals are integrated into the language model, MLLMs can be broadly categorized into “cross-modal attention” ones and “prefix-tuning” ones [80]. The former injects visual information into different layers of LLMs using cross-modal attention [4, 40], whereas the latter views the visual tokens as part of the language token sequence and directly append them with text embeddings [54, 44, 43]. Our model belongs to the prefix-tuning family by following a LLaVA-styled multimodal architecture. Considering that MLLM is a fast growing field, readers are recommended to refer to more detailed studies and surveys such as [80, 42, 23].
4.2 Vision encoder design for MLLMs
Our work is closely related to research focusing on improved vision encoder designs for MLLM. Early works [54, 44] usually adopted vision encoder pre-trained on vision-language alignment tasks such as CLIP [63] and EVA-CLIP [70]. Stronger vision encoders such as SigLIP [83] and InternVL [13] have been proposed to improve vision-language tasks with improved designs, larger model sizes, and better training recipes. Considering that the models are often pre-trained on low-resolution images and lack the ability to encode fine-grained details, higher resolution adaptation is often performed to increase the MLLM input resolution [9, 10, 6, 12].
Besides higher resolution adaptation, models such as LLaVA-NeXT [53], LLaVA-UHD [79], Monkey [47], InternLM-XComposer [17], and InternVL [12] use either tiling or adaptive tiling to handle high resolution input, where input images are divided into lower-resolution patches and process them separately. Although the ability to handle higher resolution is similarly made possible with the introduction of additional vision experts, its spirit is slightly orthogonal to the tiling techniques in terms of introducing the diversity of visual representations. Both techniques are compatible and can be combined together.
Our work is most related to existing models using multiple vision encoders for improved perception. Mini-Gemini [45] and LLaVA-HR [59] propose to fuse high-resolution visual features into the low-resolution visual tokens. Apart from the resolution issue, these pre-trained vision encoders may lack specific abilities such as reading text and localizing objects. Hence, a series of works have integrated vision models pre-trained on various vision tasks to extend the capabilities of vision encoders. For example, Mousi [20], and Brave [33] fuse visual tokens from different vision encoders by concatenating along the channel or token direction. RADIO [64] proposes a multi-teacher distillation method to unify the abilities of different vision encoders into a unified model. MoAI [37], IVE [28], and Prismer [55] further use the output of vision experts, such as OCR, detection, or depth estimation, to supplement more information for MLLM to generate answers. MoVA [88] devises a routing network to assign an optimal vision model conditioned on the given image and instructions.
5 Conclusion
We conduct an in-depth analysis study on the design space for integrating vision encoders for multimodal large language models. Unlike previous works that focus on designing novel fusion paradigms, we find systematic design choice matters and discover a series of useful techniques. Step by step, we optimize the training recipe of individual vision encoders, identify an extendable and efficient fusing method, and gradually combine vision encoders with different domain knowledge. The results show the importance of basic design space. We hope our work can serve as a new basis and bring new inspiration for the vision encoder design for MLLM.

### v2

@@ 4 Related work
5 Conclusion
We conduct an in-depth analysis study on the design space for integrating vision encoders for multimodal large language models. Unlike previous works that focus on designing novel fusion paradigms, we find systematic design choice matters and discover a series of useful techniques. Step by step, we optimize the training recipes of individual vision encoders, identify an extendable and efficient fusion method, and gradually combine vision encoders with different domain knowledge. The results show the importance of basic design space. We hope our work can serve as a new basis and bring new inspiration for the vision encoder design for MLLM.
6 Acknowledgments
The team would like to thank Yunhao Fang and Jason Lu for sharing the data and training recipes, and Yawen Luo for the assistance on figure editing. We thank Wei Ping, Zhuolin Yang, Wenliang Dai, Nayeon Lee, Boxin Wang, Ilia Karmanov, Lukas Voegtle, Philipp Fischer, Matthieu Le and Tuomas Rintamaki for their assistance on the internal codebases. We also thank the valuable discussions and input from Zhiqi Li, Guo Chen, Shilong Liu, Jihao Liu, Ming-Chang Chiu, Yunze Man, Shiyi Lan, Nadine Chang, Maying Shen, Vibashan VS, Jenny Schmalfuss, Jose Alvarez, Amala Sanjay Deshmukh, Mike Ranzinger, Greg Heinrich, Pavlo Molchanov, Vidya Murali, Parthasarathy Sriram, Mohammad Shoeybi, Song Han, Ofri Masad, Osvald Nitski, Qing Miao, Yao Xu, Jane Scowcroft, Dmitry Chichkov and Padmavathy Subramanian. Finally, the team would like to thank the Hugging Face Team that for the support of ZERO GPU demo, and the NVIDIA infrastructure team for their prompt and helpful assistance. Min Shi is partly supported by NSF Award #2427478 and #2229873.
Appendix A Appendix
A.1 Benchmark details
In this section, we provide the additional benchmark details for the tables in Section 2. The detailed comparison of different adaptation methods for the CLIP encoder are shown in Table 9. Table 10 shows the comparison between different vision encoders on all the adopted benchmarks. Table 11 list the detailed results on the vision encoder fusion methods. Table 12 shows the comparison between different vision encoder combinations on each benchmark.
Table 9: Comparison of different high-resolution adaption methods to strengthen CLIP model (336x336). RWQA denotes the RealworldQA (xAI, 2024).
Vision Encoder Unfreeze Res GQA VizWiz MME SEED OCR DocVQA ChartQA AI2D POPE RWQA SQA Avg.
Original 336 63.2 55.1 1574 70.2 354 59.6 42.1 71.1 86.7 58.0 72.5 616.5
Original 336 60.9 54.0 1501 61.8 305 56.4 24.5 69.3 80.9 53.6 72.0 562.6
Interpolate 448 62.7 53.8 1470 69.1 299 58.4 37.7 71.2 85.7 55.3 69.6 589.7
Interpolate 448 65.6 57.8 1534 73.7 526 65.0 61.1 73.7 87.3 57.7 71.5 670.5
Interpolate 672 64.9 55.7 1503 72.4 509 64.6 62.0 72.2 87.1 57.4 71.2 674.2
Tiled-input 672 63.0 54.9 1529 72.5 435 64.9 65.7 71.5 87.6 57.0 71.4 673.9
InternVL 448 63.6 56.9 1537 71.7 529 65.0 58.6 72.9 87.4 59.2 70.2 661.9
InternVL 448 65.6 57.8 1534 73.7 526 65.0 61.1 73.7 87.3 58.8 71.5 671.5
Table 10: Comparison between different vision experts as the MLLM encoders.
Category Vision Encoder Unfreeze Res GQA VizWiz MME SEED OCR DocVQA ChartQA AI2D POPE RWQA SQA Avg.
VL Alignment ConvNeXt 1024 63.3 53.5 1526 70.6 404 70.4 60.8 71.6 87.5 57.1 68.6 635.0
1024 63.3 54.4 1510 72.9 518 77.9 67.0 72.1 88.1 58.8 68.6 659.7
Segmentation SAM 1024 57.3 49.0 1216 56.9 38 20.1 17.4 69.2 84.3 49.2 66.8 471.1
1024 60.2 51.5 1291 65.9 35 21.2 17.8 70.7 86.4 54.1 65.7 494.7
Object Detection EVA-02 1024 63.1 51.1 1359 69.2 123 25.6 25.2 71.2 88.5 57.9 66.1 523.5
1024 64.3 55.5 1449 72.7 358 57.1 57.5 72.2 88.3 59.6 67.7 614.4
Text Recognition Pix2Struct 1024 53.1 48.1 1296 53.4 460 71.0 61.0 69.6 79.2 46.7 65.5 578.7
1024 54.9 47.3 1262 55.1 472 72.5 62.0 68.7 80.0 49.3 66.6 584.8
Self-Supervised DINOv2 448 62.4 53.1 1438 67.4 41 20.2 17.3 70.7 85.3 53.3 67.1 503.1
448 64.2 55.5 1466 71.8 45 20.4 17.5 71.4 87.4 57.3 67.8 518.0
Table 11: Comparison of different fusion methods for different vision experts. “#Tokens(V)” denotes the number of visual tokens.
Vision Encoders Fusion #Tokens(V) GQA VizWiz MME SEED OCR DocVQA ChartQA AI2D POPE RWQA SQA Avg.
CLIP + ConvNeXt Seq. Append 2048 64.8 54.5 1563 73.4 532 77.7 67.6 72.4 87.9 61.2 68.8 690.5
Channel Concat. 1024 63.2 48.0 1497 73.5 551 77.7 67.0 72.4 88.3 59.1 70.7 681.5
LLaVA-HR 1024 64.5 57.2 1538 72.0 498 74.5 63.8 72.3 87.7 59.2 68.7 678.7
Mini-Gemini 1024 65.3 56.9 1548 72.9 478 68.3 63.2 71.5 87.3 59.7 69.4 672.5
Deformable Attn. 1024 64.0 57.3 1504 72.7 463 69.5 64.4 73.3 87.4 62.8 68.9 674.3
CLIP + ConvNeXt + SAM Seq. Append 3072 64.3 53.6 1539 73.2 525 77.9 67.0 72.3 87.4 60.1 69.5 686.2
Channel Concat. 1024 63.3 55.9 1528 73.3 545 78.9 67.2 72.3 88.4 59.2 70.0 690.4
Table 12: Detailed comparison on vision encoder combinations.
#Encoder Encoder Combination GQA VizWiz MME SEED OCR DocVQA ChartQA AI2D POPE RWQA SQA Avg.
2 CL + CN 63.2 48.0 1497.0 73.5 551.0 77.7 67.0 72.4 88.3 59.1 70.7 681.5
3 CL + CN + DI 63.3 55.9 1528.0 73.3 545.0 78.9 67.2 72.3 88.4 59.2 70.0 690.4
CL + CN + SA 64.6 55.3 1504.0 73.3 526.0 75.7 64.9 72.1 88.3 61.1 70.9 685.4
CL + CN + PS 63.2 51.4 1497.0 73.3 550.0 78.5 65.9 73.1 87.7 60.3 70.5 685.1
CL + CN + EV 63.2 51.7 1565.0 73.9 538.0 77.7 67.8 73.6 89.0 61.4 69.4 690.7
CL + CN + EV + DI 63.6 54.9 1512.0 73.8 547.0 77.0 66.7 73.1 88.9 60.4 69.7 689.4
CL + CN + EV + SA 64.3 57.7 1533.0 73.7 521.0 75.2 65.3 72.2 88.5 61.1 70.2 688.0
CL + CN + EV + PS 64.8 56.5 1561.0 73.4 540.0 78.8 67.5 72.2 88.4 59.9 70.5 694.6
5 CL + CN + EV + PS + SA 64.7 59.1 1528.0 73.9 529.0 78.6 67.8 72.9 88.8 62.2 69.7 697.1
CL + CN + EV + PS + DI 64.7 54.1 1506.0 73.7 541.0 75.1 64.9 72.7 88.3 60.0 70.3 684.7
6 CL + CN + EV + PS + SA + DI 63.8 57.8 1512.0 73.5 525.0 75.1 65.8 71.8 88.4 61.4 69.9 686.8
Table 13: Comparison between different training strategies. ‘1 epoch” means we train Eagle for 1 epoch in the supervised fine-tuning stage. ‘unlock*” means we unlock vision encoders in the pre-training stage.
Config Prealign Pretrain Finetune
GQA
MME
OCR
SciQA
POPE
DocVQA
ChartQA
SEED
Vizwiz
AI2D
RWQA
Avg.
1 epoch llava595k Eagle1.8M 64.7 1528 52.9 69.7 88.8 78.6 67.7 73.9 59.1 72.8 62.2 697.1
2 epoch llava595k Eagle1.8M 65.4 1539 51.4 70.3 87.9 79.8 67.9 73.8 58.5 73.5 62.7 698.3
1 epoch, unlock* llava595k Eagle1.8M 64.1 1541 54.4 71.5 88.5 79.1 68.5 74.0 56.6 72.2 61.9 698.0
1 epoch, unlock* llava595k+Eagle1.8M Eagle1.8M 65.3 1545 54.8 70.5 88.5 78.8 68.4 73.5 57.5 72.0 62.9 699.5
1 epoch Eagle1.8M llava595k Eagle1.8M 64.5 1579 55.4 71.6 88.7 79.3 68.9 73.9 58.9 73.9 63.4 706.6
1 epoch, unlock* Eagle1.8M llava595k Eagle1.8M 65.1 1561 56.0 71.3 88.5 79.3 69.5 74.2 59 73.5 63.4 707.1
1 epoch, unlock* llava595k+Eagle1.8M llava595k Eagle1.8M 64.5 1570 56.0 71.7 89 79.7 68.9 74.0 58.5 74.0 63.7 707.8
1 epoch, unlock* llava595k+Eagle1.8M llava595k+Eagle1.8M Eagle1.8M 64.6 1582 56.6 71.9 88.7 79.8 69.3 73.7 58.7 73.5 63.9 708.9
A.2 Additional qualitative results
Fig. 6 demonstrates that our model can perform reasoning based on the knowledge to answer some difficult questions. Fig. 7 shows some examples of OCR and document understanding cases.